# GigaSpeech 2: An Evolving, Large-Scale and Multi-domain ASR Corpus for Low-Resource Languages with Automated Crawling, Transcription and Refinement

### v1

@@ B.1 Domain Distribution of Manual Evaluation Sets
Figure 2: Hours distribution of manual evaluation sets for Thai, Indonesian and Vietnamese. The inner circle represents the format, and the outer circle represents the topic.
B.2 Duration Distribution of Training Sets
@@
(a) Thai raw
(b) Thai refined
(c) Indonesian raw
(d) Indonesian refined
(e) Vietnamese raw
(f) Vietnamese refined
Figure 3: Utterance-level duration (second) distribution of training sets for Thai, Indonesian and Vietnamese.
Appendix C Model Configurations
C.1 Configuration of Zipformer
Two Zipformer-based models are used, following official configurations reported in icefall [^8] https://github.com/k2-fsa/icefall. In each Zipformer stack, the hidden dimensions of the first and last feedforward modules are 3/4 and 5/4 of the middle one, respectively. Ahead of the encoder, a convolution subsampling module with a stride of 2 reduces the frame rate to 50 Hz. The input consists of 80-channel FBank features extracted over windows of 25ms, strided by 10ms. The label decoder utilizes a stateless decoder [15]. 8 V100 32G GPUs are used for training. Detailed configurations are provided in Table 6.
@@ C.1 Configuration of Zipformer
C.2 Configuration of Conformer
@@
Appendix D Ablation Study on Noisy Student Training
Table 8: Ablation study of NST on GigaSpeech 2 Thai, evaluated across various evaluation sets: GigaSpeech 2 DEV and TEST, Common Voice 17.0 TEST, and FLEURS TEST.
NST method # Hours (h) CER
GigaSpeech 2 DEV   TEST Common Voice TEST FLEURS TEST
Sys. 1 (Tab. 2, iter 2 $\rightarrow$ iter 3) 7219 10.47 12.38 4.63 10.96
Sys. 2 (Tab. 2, iter 2 $\rightarrow$ iter3, without relabeling) 7219 10.77+2.9% 12.90+4.2% 5.23+13.0% 10.72-2.2%
Sys. 3 (Tab. 2, iter 2 $\rightarrow$ iter 3, larger augmentation) 7219 10.65+1.7% 12.81+3.5% 5.36+15.8% 10.86-0.9%
Based on the ablation study of our proposed NST on the evaluation sets in Table 8, we can analyze the effects of different iterations and their impact on performance: 1) Relabeling the data during the transition from iteration 2 to 3 is crucial for improving performance (Sys.1 vs. Sys.2). 2) Larger augmentation applied in our NST process may have a negative impact on the performance (Sys.1 vs. Sys.3). These findings suggest that careful consideration of the relabeling and augmentation strategies is crucial for optimizing the performance of the NST model across different evaluation sets and domains.

### v2

@@ A.2 Domain Distribution of Manual Evaluation Sets
Figure 2: Hours distribution of manual evaluation sets for Thai (th), Indonesian (id), and Vietnamese (vi). The inner circle represents the format, and the outer circle represents the topic.
A.3 Duration Distribution of Training Sets
@@
(a) th raw
(b) th refined
(c) id raw
(d) id refined
(e) vi raw
(f) vi refined
Figure 3: Utterance-level duration (second) distribution of training sets for Thai (th), Indonesian (id), and Vietnamese (vi).
A.4 Evaluation of Processing Time
The processing times for transcription, forced alignment, filtering, segmentation, and relabeling are measured on an idle single V100 32G GPU machine using a 100-hour subset of Thai audio. The processing time and the real-time factor (RTF) are detailed in Table 6.
Table 6: Evaluation of overall processing time and real-time factor (RTF) for each process in the construction of GigaSpeech 2. The processing times for transcription, forced alignment, filtering, segmentation, and relabeling are measured on an idle single V100 32G GPU machine using a 100-hour subset of Thai audio.
Process Time Consumption RTF
Transcription 19h 42min 13s $1.97\times 10^{-1}$
Forced Alignment 3h 27min 29s $3.46\times 10^{-2}$
Filter 3s $8.00\times 10^{-6}$
Segmentation 6min 58s $1.16\times 10^{-3}$
Relabel 40min 48s $6.80\times 10^{-3}$
Appendix B Model Configurations
@@
Table 8: Ablation study of NST on GigaSpeech 2 Thai, evaluated across various evaluation sets, including GigaSpeech 2 DEV and TEST, Common Voice 17.0 TEST, and FLEURS TEST.
NST method CER
GigaSpeech 2 DEV   TEST Common Voice TEST FLEURS TEST
Sys. 1 (Tab. 2, iter 2 $\rightarrow$ iter 3) 10.47 12.38 4.63 10.96
Sys. 2 (Tab. 2, iter 2 $\rightarrow$ iter 3, without relabeling) 10.77+2.9% 12.90+4.2% 5.23+13.0% 10.72-2.2%
Sys. 3 (Tab. 2, iter 2 $\rightarrow$ iter 3, larger augmentation) 10.65+1.7% 12.81+3.5% 5.36+15.8% 10.86-0.9%
@@
B.1 Configuration of Zipformer
Two Zipformer-based models are used, following official configurations reported in icefall [^11] https://github.com/k2-fsa/icefall. In each Zipformer stack, the hidden dimensions of the first and last feedforward modules are 3/4 and 5/4 of the middle one, respectively. Ahead of the encoder, a convolution subsampling module with a stride of 2 reduces the frame rate to 50 Hz. The input consists of 80-channel FBank features extracted over windows of 25ms, strided by 10ms. The label decoder utilizes a stateless decoder Ghodsi et al. (2020). 8 V100 32G GPUs are used for training. Detailed configurations are provided in Table 7.
B.2 Configuration of Conformer
Appendix C Ablation Study on Noisy Student Training
Based on the ablation study of our proposed NST on the evaluation sets in Table 8, we can analyze the effects of different iterations and their impact on performance:
1) Relabeling the data during the transition from iteration 2 to 3 is crucial for improving performance (Sys.1 vs. Sys.2).
2) Larger augmentation applied in our NST process may hurt the performance (Sys.1 vs. Sys.3).
These findings suggest that careful consideration of the relabeling and augmentation strategies is crucial for optimizing the performance of the NST model across different evaluation sets and domains.
Table 10: ASR performance of Whisper Medium with/without fine-tuning on GigaSpeech 2 Thai, tested on GigaSpeech 2 TEST and Common Voice 17.0 TEST
Model CER
GigaSpeech 2 Common Voice
Whisper medium 37.55 16.41
+ GigaSpeech 2 Thai fine-tuned 14.15-62.3% 6.92-57.8%
Appendix D Additional Results of Whisper Medium
We evaluated Whisper medium [^13] https://huggingface.co/openai/whisper-medium and its fine-tuned version [^14] https://huggingface.co/scb10x/monsoon-whisper-medium-gigaspeech2 on GigaSpeech 2 Thai, using the test sets from GigaSpeech 2 and Common Voice 17.0. As shown in Table 10, fine-tuning resulted in an approximate 60% relative CER reduction across two test sets, indicating the high quality of the GigaSpeech 2 Thai.